# AnimateAnywhere: Rouse the Background in Human Image Animation https://animateanywhere.github.io/

### v1

@@ 4.2. Comparison with Existing Methods
We compare our method with existing state-of-the-art human animation methods in both quantitative and qualitative aspects, including MagicAnimate (Xu et al., 2024b), AnimateAnyone (Hu, 2024), Champ (Zhu et al., 2024), MusePose (Tong et al., 2024), and HumanVid (Wang et al., 2024a). Since AnimateAnyone (Hu, 2024) is not officially open-sourced, we adopt its publicly available reimplementation from Moore-AnimateAnyone111https://github.com/MooreThreads/Moore-AnimateAnyone for evaluation. Note that these methods, except HumanVid (Wang et al., 2024a), are originally trained on datasets with static backgrounds. For fair comparison, we retrain AnimateAnyone (Hu, 2024) and MusePose (Tong et al., 2024) on our training dataset. In addition, Humanvid (Wang et al., 2024a) is capable of generating dynamic backgrounds, similar to ours. As its official implementation is not released, we reimplement Humanvid (Wang et al., 2024a) according to the details mentioned in its paper.
Quantitative Comparison. We conduct comprehensive experiments to assess the effectiveness of the proposed method, and the quantitative results are shown in table 1. From all metrics on the Humanvid (Wang et al., 2024a) test dataset and the BL200 test dataset, we can see that previous methods struggle to generate videos with dynamic backgrounds, resulting in the difference between the dynamic ground-truth videos. Even the retrained versions of these baselines occasionally produce static backgrounds, as they are not properly designed for background motion synchronized with human poses. In contrast, our method leverages CogVideoX (Yang et al., 2024b), the BML, and the epipolar loss to achieve substantial improvements, outperforming all baselines in reconstruction metrics. Furthermore, it also achieves state-of-the-art video fidelity, improving FID by 6% and FVD by 12% over the strongest baseline.
@@
Qualitative Comparison. The qualitative comparison with competing methods is shown in fig. 4. MagicAnimate (Xu et al., 2024b), AnimateAnyone (Hu, 2024), Champ (Zhu et al., 2024), and MusePose (Tong et al., 2024) generate videos with static backgrounds that are usually unrealistic and exhibit various artifacts. Even after retraining, these methods show limited improvements and still fail in some circumstances. For instance, the left example exhibits the static background, while the right example produces an inharmonious scene in which only the wall moves and the table remains motionless. Humanvid is more likely to generate dynamic backgrounds, yet the background motion is frequently misaligned with the human. In the left example, the background moves noticeably slower than the person. In contrast, our method synthesizes photorealistic videos while maintaining coherent motion between the human and the background. fig. 5 illustrates more generated videos on BL200 dataset. These results also highlight our method’s capability to generate realistic background motion while maintaining high-quality, temporally consistent animation of both the human and the background. Moreover, our method achieves better identity-preserving than other methods.
PSNR | SSIM | LIPIS | FID | FVD | |
Ours w/o | 18.06 | 0.5807 | 0.3246 | 45.22 | 981.37 |
Ours with () | 17.32 | 0.5697 | 0.3280 | 50.81 | 1070.79 |
Ours with ()) | 18.07 | 0.5831 | 0.3119 | 44.51 | 978.09 |
@@
5. Ablation Study
@@
Effects of Epipolar Loss. As shown in fig. 7, the generated video exhibits noticeable artifacts in the background without incorporating epipolar loss, because the learned background motion does not adhere to geometric constraints. Additionally, suppressing all attention values outside the epipolar mask does not alleviate this issue but performs even worse. As evidenced in the second row of fig. 7, noticeable artifacts and unexpected objects appear in the background. In contrast, our approach achieves desired results by applying to the region of . As shown in table 2, our method with outperforms both the one without epipolar loss and the one using the mask , demonstrating its effectiveness.
@@
Effects of VGG Loss. We investigate the impact of the VGG perceptual loss on video generation. While the 3D VAE used in CogVideoX (Yang et al., 2024b) achieves efficient spatiotemporal compression (8 spatially, 4 temporally), we observe that optimizing solely with an L2 loss in latent space leads to degraded high-frequency details. As shown in fig. 8, the hand regions exhibit blurred contours without . After incorporating in pixel space, the details of the hand regions are well-preserved.
Effects of Foundational Models. To validate the superiority of our base model CogVideoX (Yang et al., 2024b), we select several state-of-the-art video generation base models (AnimateDiff (Guo et al., 2023), SVD (Blattmann et al., 2023)) for comparison. Specifically, we build two new pipelines based on AnimateDiff (Wang et al., 2024a) and SVD (Blattmann et al., 2023), and train them using our training dataset. Although these pipelines can generate videos with dynamic backgrounds, they struggle with the generation of novel regions. For instance, as shown in fig. 9, the pipeline based on AnimateDiff (Guo et al., 2023) generates blurred novel regions, while the pipeline based on SVD (Blattmann et al., 2023) yields unnatural human parts. In contrast, our method, based on CogvideoX (Yang et al., 2024b), successfully addresses these challenges.
6. Conclusion

### v2

@@ IV-B Comparison with Existing Methods
We compare our method with existing state-of-the-art human animation methods in both quantitative and qualitative aspects, including MagicAnimate [2], AnimateAnyone [1], Champ [3], MusePose [19], DynamiCtrl [51], and Humanvid [4]. Since AnimateAnyone [1] is not officially open-sourced, we adopt its publicly available reimplementation from Moore-AnimateAnyone111https://github.com/MooreThreads/Moore-AnimateAnyone for evaluation. Note that these methods, except Humanvid [4], are originally trained on datasets with static backgrounds. For fair comparison, we retrain AnimateAnyone [1] MusePose [19] and DynamiCtrl [51] on our training dataset. Specifically, AnimateAnyone [1] and MusePose [19] are built upon the AnimateDiff [27], while DynamiCtrl [51] adopts CogVideoX [9] as base models.
Quantitative Comparison. We conduct comprehensive experiments to assess the effectiveness of the proposed method, and the quantitative results are shown in table I. From all metrics on the Humanvid [4] test dataset and the BL200 test dataset, we can see that previous methods originally struggled to generate videos with dynamic backgrounds, resulting in significantly lower metrics. Even the retrained versions of AnimateAnyone [1], MusePose [19], and DynamiCtrl [51] still yield inferior results as they still exhibit static or inconsistent backgrounds. Humanvid [4] achieves reasonable performance by relying on precisely estimated camera poses. In contrast, our method does not depend on camera pose and instead leverages the BML module together with the epipolar loss to effectively model dynamic backgrounds. As a result, our approach not only exhibits significant improvement over baseline methods but also surpasses Humanvid [4] on most metrics.
@@
Qualitative Comparison. The qualitative comparison with competing methods is shown in fig. 4. MagicAnimate [2], AnimateAnyone [1], Champ [3], MusePose [19], and DynamiCtrl [51] generate videos with static backgrounds that are usually unrealistic. Even after retraining, AnimateAnyone [1] and MusePose [19] show only marginal improvements and still fail to generate dynamic backgrounds in many cases. For example, the left case in fig. 4 exhibits a completely static background, while the right case yields an inconsistent scene where only the wall moves, but the table remains still. Although retrained DynamiCtrl [3] and Humanvid are more capable of dynamic backgrounds, the generation regions are often inconsistent with the reference background or contain noticeable artifacts, which are highlighted by red boxes in fig. 4. In contrast, our method synthesizes photorealistic videos while maintaining coherent motion between the human and the background. We also provide more generated videos on the BL200 dataset in fig. 5. These results also highlight our method’s capability to generate realistic background motion while maintaining high-quality, temporally consistent animation of both the human and the background. Moreover, our method achieves superior identity preservation, particularly in face regions, compared with other methods. We provide videos of these examples and more examples on the project page.
| Method | PSNR | SSIM | LPIPS | FID | FVD |
| Ours w/o | 15.83 | 0.4823 | 0.4063 | 96.48 | 1269.42 |
| Ours with () | 14.71 | 0.4315 | 0.4261 | 114.26 | 1373.42 |
| Ours with ()) | 16.01 | 0.5831 | 0.3886 | 91.29 | 1243.07 |
@@
IV-C Ablation Study
| Method | PSNR | SSIM | LPIPS | FID | FVD |
| Explicit Two-stage | 17.00 | 0.5646 | 0.3841 | 61.07 | 1604.10 |
| Ours w/o BML | 16.86 | 0.5676 | 0.3333 | 51.15 | 1199.54 |
| Ours | 18.07 | 0.5831 | 0.3119 | 44.51 | 978.09 |
Effects of Background Motion Learner Module. As illustrated in section III-B, we introduce a Background Motion Learner (BML) module that directly infers background motion from human pose sequences. In this subsection, we evaluate three strategies for background motion learning: (1) an explicit two-stage approach, (2) without BML, and (3) the proposed BML module. As shown in table III, the two-stage method suffers from error accumulation and scale inconsistency in labeled trajectories, leading to degraded performance. In contrast, our BML reformulates the task as an end-to-end background motion learning problem, achieving the best overall results. Furthermore, compared with the model without BML, our method improves PSNR by 7.2% and FID by 13.0%, clearly validating the effectiveness of the proposed module.
@@ IV-C Ablation Study
Effects of Epipolar Loss. In section III-C, we designed the epipolar loss to better learn background motion. However, some videos in Humanvid [4] test dataset contain static or weak background motion, which diminishes its effect. To better assess the effects of epipolar loss, we select a subset with significant motion based on camera parameters and of the start and end frames: where and are the 10th largest translation and rotation differences in the test set. As shown in table II, our method with outperforms both the one without epipolar loss and the one using the mask , demonstrating its effectiveness. Qualitative comparisons in fig. 7 further illustrate this. Without epipolar loss, the generated backgrounds exhibit noticeable artifacts as the learned background motion does not adhere to geometric constraints. Additionally, suppressing all attention values outside the epipolar mask does not alleviate this issue but performs even worse. As evidenced in the second row of fig. 7, noticeable artifacts and unexpected objects appear in the background. In contrast, our approach achieves desired results by applying to the region of .
@@
Effects of VGG Loss. We investigate the impact of the VGG perceptual loss on video generation. While the 3D VAE used in CogVideoX [9] achieves efficient spatiotemporal compression (8 spatially, 4 temporally), we observe that optimizing solely with an L2 loss in latent space leads to degraded high-frequency details. As shown in fig. 8, the hand regions exhibit blurred contours without . After incorporating in pixel space, the details of the hand regions are well-preserved.
| Value | PSNR | SSIM | LPIPS | FID | FVD |
| 10 | 18.05 | 0.5804 | 0.3131 | 45.64 | 984.23 |
| 30 | 18.07 | 0.5831 | 0.3119 | 44.51 | 978.09 |
| 50 | 17.02 | 0.5671 | 0.3217 | 46.72 | 1087.21 |
| 70 | 16.87 | 0.5571 | 0.3375 | 47.77 | 1229.87 |
| 90 | 16.85 | 0.5479 | 0.3455 | 50.51 | 1289.06 |
Effects of threshold . In our adaptive epipolar constraint, the threshold is used to identify low-confidence attention regions. We perform a sensitivity analysis by varying from the 0th to the 100th percentile, as shown in table IV. When the threshold is too small (), the epipolar constraint is inactive. In contrast, a high threshold () may interfere with the learned 3D attention, leading to degraded performance. We observe that setting to the 30th percentile yields the best performance and thus adopt this value.
V Limitations.
Although AnimateAnywhere performs well in generating human animations with vivid and realistic backgrounds, it encounters challenges in predicting background motion under certain ambiguous conditions. In such cases, a single human pose sequence may correspond to multiple plausible background motions. For example, considering a boy standing on a skateboard and appearing to grow larger, it is unclear whether the boy is skateboarding forward with the camera slowly zooming out (the boy’s forward motion is faster than the camera’s zoom-out) or the boy is standing still with the camera zooming in. This ambiguity arises because the action of skateboarding is difficult to infer solely from the human pose sequence. Nevertheless, regardless of the above scenario, our method generates videos with plausible background motion as shown in fig. 9. We provide corresponding videos on the project page. To address this issue in the future, we plan to incorporate text descriptions that explicitly specify the intended background motion. This can help reduce ambiguity and better align the generated video with user expectations.
VI Conclusion